# A Clique Partitioning-Based Algorithm for Graph Compression

### v2

@@ A Appendix
Table 1: Degrees of vertices in partition $W$ at different steps of CPGC.
@@
Figure 6: Clique partitioning: (a) given bipartite graph $G(U,W)$; (b) extracted clique $C_{1}$; (c) extracted clique $C_{2}$; (d) graph with trivial cliques; and (e) compressed tripartite graph $G(U,Z,W)$
@@
In Line 5, CSA forms set $\mathcal{K}$, where $\mathcal{K}=\{w_{4},w_{2},w_{3},w_{5},w_{6}\}$ has all vertices whose degrees are greater than or equal to $d_{w_{\pi(\hat{k})}}=7$. This results in $\gamma=2$ in Line 7, therefore CSA extracts two cliques in Lines 8-13. Since $q$ is initialized to 0 in CPGC, the for loop in Line 8 iterates two times for $c=1$ and $2$. In Line 9, CSA forms $K_{1}=\{w_{4},w_{2}\}$, and in Line 10, it forms the left partition $U_{K_{1}}$ of bipartite clique $\mathcal{C}_{1}$, by selecting the common neighbors of the vertices in $K_{1}$. For example, for set $K_{1}=\{w_{4},w_{2}\}$, $U_{K_{1}}=U-\{u_{7}\}$ as $w_{4}$ and $w_{2}$ have $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{6},u_{8}\}$ as common neighbors. In Line 12, CSA updates the adjacency matrix $A$ by removing the edges of bipartite clique $\mathcal{C}_{1}$ from the original bipartite graph $G$ and finally in Line 13, it updates all $d_{w}$ by subtracting $|U_{K_{1}}|=7$ from both $d_{w_{4}}$ and $d_{w_{2}}$, thus $d_{w_{4}}=1$ and $d_{w_{2}}=0$. Similarly, when $c=2$, $K_{2}=\{w_{3},w_{5}\}$ and $U_{K_{2}}=U-\{u_{6}\}$. CSA updates the adjacency matrix $A$ and $d_{w}$ such that $d_{w_{3}}=d_{w_{5}}=0$. Therefore, in the first execution, CSA, forms two bipartite cliques and removes a total of $(7\times 2)+(7\times 2)=28$ edges from $G$. The updated degrees of the vertices are shown in Table 1. It then forms the set of bipartite cliques extracted in the current execution in Line 14, $\mathscr{C}=\{C_{1},C_{2}\}$. The two cliques are shown in Figures 6(b) and 6(c).
@@
CSA returns $(\mathscr{C},\hat{A},\hat{d}_{w})$ and then CPGC updates the set $\mathscr{C}=\{C_{1},C_{2}\}$, $d_{w}$, the adjacency matrix $A$, and the number of edges $\hat{m}=54-28=26$, in Lines 8-10. In Line 12, CPGC updates $q=2$ and finally, in Line 13 it updates $\hat{k}$ according to the new value of $\hat{m}$ in Line 11 which results in $\hat{k}=1$. Since $\hat{k}=1$, it means CPGC extracts only trivial bipartite cliques (shown in Figure 6(d)) which do not contribute to graph compression. Thus, it does not meet the condition in the while loop (Line 6) and CPGC terminates.
@@
CPGC then compresses the graph by adding two vertices, $z_{1}$, and $z_{2}$, corresponding to the two cliques $C_{1}$ and $C_{2}$, and adds the corresponding edges to form the tripartite graph. The edges in the given graph $G(U,W,E)$, that are not part of any $\delta$-clique are connected directly as shown in Figure 6(e).
A.3 Proofs
Proof of Theorem 4.1.
  • Proof.
    Let’s assume that CPGC extracts only one $\delta$-clique $C_{q}$ from the given graph $G$. In the $\delta$-clique $C_{q}$, the right partition $K_{q}\subseteq W$ is formed by selecting $\hat{k}$ vertices from $W$ and the left partition $U_{q}\subseteq U$ is formed with the common neighbors of $K_{q}$, forming a complete bipartite graph with $E_{q}$ edges. The compressed graph $G^{*}$ is formed with the same left and right partitions $U$ and $W$ of the graph $G$, and a third partition $Z$ which is a set of additional vertices associated with each of the $\delta$-cliques that CPGC extracts. Our main concern is with the edges $(u_{i},w_{j})\in E^{*}$ in the compressed graph $G^{*}$. $E^{*}$ contains the edges that replace the $\delta$-clique $C_{q}$ by adding the additional vertex $z_{q}\in Z$, and the set of edges $\hat{E}=E\setminus E_{q}$ which were not part of the $\delta$-clique $C_{q}$. It can be easily seen that $\hat{E}\subset E$, that is, the edges in $\hat{E}$ are edges in $G$, and the remaining edges $E^{*}\setminus\hat{E}$ in $E^{*}$ are the edges that replace $C_{q}$ in $G^{*}$. Each edge $(u_{i},w_{j})\in E_{q}$, is replaced by two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, where $u_{i}\in U$, $w_{j}\in W$, and $z_{q}\in Z$. Therefore, for each edge $(u_{i},w_{j})\in E_{q}$, there exists a path from $u_{i}$ to $w_{j}$ composed of two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, that passes through the additional vertex $z_{q}\in Z$, thus preserving the path information. This holds true for all the $\delta$-cliques extracted by CPGC.
Proof of Lemma 4.1.
  • Proof.
    In the $\delta$-clique based graph compression performed by CPGC, the compressed graph $G^{*}(U,W,Z,E^{*})$ is obtained from $G$ by replacing $|E_{q}|=|U_{q}\times K_{q}|$ edges in a $\delta$-clique $C_{q}$ with $|E^{*}|=|U_{q}|+|K_{q}|$ edges in $G^{*}$ and adding an additional vertex $z_{q}\in Z$ for each extracted $\delta$-clique $C_{q}$. The size of the right partition of $C_{q}$ is $|K_{q}|={k}=k(n,m,\delta)=\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/m)}}\Big{\rfloor}$. If $m<2n^{2-\frac{\delta}{2}}$ then the size of the right partition $|K_{q}|$ is $k=1$. Therefore, the number of edges in the $\delta$-clique is $|E_{q}|=|U_{q}|\cdot 1=|U_{q}|$. Those edges are replaced by $|E^{*}|=|U_{q}|+\hat{k}=|U_{q}|+1$ edges in $G^{*}$. Thus, replacing the $\delta$-clique in $G$ with $m<2n^{2-\frac{\delta}{2}}$, the number of edges in $G^{*}$ would actually increase by 1, which does not lead to a compression of $G$.
A.4 Extension to non-bipartite graphs.
CPGC can be extended to compress non-bipartite graphs using a similar technique to that described in [9]. Consider a directed graph $H(V,E)$ with $n$ vertices and $m$ edges. Initially, we transform $H$ into a bipartite graph $G\left(L,R,E^{\prime}\right)$, where each vertex $v\in V$ is duplicated into $v_{L}\in L$ and $v_{R}\in R$. For each directed edge $(u,v)$ in $E$, we create a directed edge $\left(u_{L},v_{R}\right)$ in $E^{\prime}$, with the direction from $L$ to $R$. Furthermore, for each vertex $v\in V$ we add a directed edge from $v_{R}$ to $v_{L}$. The proposed compression algorithm is then applied to the bipartite graph $G$, which includes only the edges from $L$ to $R$. Once the compressed graph is computed we add to it the $n$ directed edges from $u_{R}$ to $u_{L}$, corresponding to each vertex $u\in V$. Since the path information of the original graph is preserved by this transformation, it allows different graph algorithms to work on the compressed graph as well. Undirected general graphs can be transformed first into directed graphs by simply replacing each undirected edge $(u,v)$ in the original graph by two directed edges $(u,v)$ and $(v,u)$, and then applying the technique presented above to transform the directed graph into a bipartite graph. The time complexity and the compression ratio of CPGC are the same as in the case of the bipartite graphs. A similar approach for converting a general graph to a bipartite graph and then compressing the obtained bipartite graph was employed in [3, 5].

### v3

@@ A.2 Example: CPGR Execution
Table 3: Degrees of vertices in partition $W$ at different steps of CPGR.
@@
Figure 10: Clique partitioning: (a) given bipartite graph $G(U,W)$; (b) extracted clique $C_{1}$; (c) extracted clique $C_{2}$; (d) graph with trivial cliques; and (e) restructured tripartite graph $G^{*}(U,Z,W)$
@@
In Line 5, CSA forms set $\mathcal{K}$, where $\mathcal{K}=\{w_{4},w_{2},w_{3},w_{5},w_{6}\}$ has all vertices whose degrees are greater than or equal to $d_{w_{\pi(\hat{k})}}=7$. This results in $\gamma=2$ in Line 7, therefore CSA extracts two cliques in Lines 8-13. Since $q$ is initialized to 0 in CPGR, the for loop in Line 8 iterates two times for $c=1$ and $2$. In Line 9, CSA forms $K_{1}=\{w_{4},w_{2}\}$, and in Line 10, it forms the left partition $U_{K_{1}}$ of bipartite clique $\mathcal{C}_{1}$, by selecting the common neighbors of the vertices in $K_{1}$. For example, for set $K_{1}=\{w_{4},w_{2}\}$, $U_{K_{1}}=U-\{u_{7}\}$ as $w_{4}$ and $w_{2}$ have $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{6},u_{8}\}$ as common neighbors. In Line 12, CSA updates the adjacency matrix $A$ by removing the edges of bipartite clique $\mathcal{C}_{1}$ from the original bipartite graph $G$ and finally in Line 13, it updates all $d_{w}$ by subtracting $|U_{K_{1}}|=7$ from both $d_{w_{4}}$ and $d_{w_{2}}$, thus $d_{w_{4}}=1$ and $d_{w_{2}}=0$. Similarly, when $c=2$, $K_{2}=\{w_{3},w_{5}\}$ and $U_{K_{2}}=U-\{u_{6}\}$. CSA updates the adjacency matrix $A$ and $d_{w}$ such that $d_{w_{3}}=d_{w_{5}}=0$. Therefore, in the first execution, CSA, forms two bipartite cliques and removes a total of $(7\times 2)+(7\times 2)=28$ edges from $G$. The updated degrees of the vertices are shown in Table 3. It then forms the set of bipartite cliques extracted in the current execution in Line 14, $\mathscr{C}=\{C_{1},C_{2}\}$. The two cliques are shown in Figures 10(b) and 10(c).
@@
CSA returns $(\mathscr{C},\hat{A},\hat{d}_{w})$ and then CPGR updates the set $\mathscr{C}=\{C_{1},C_{2}\}$, $d_{w}$, the adjacency matrix $A$, and the number of edges $\hat{m}=54-28=26$, in Lines 12-15. In Line 16, CPGR updates $q=2$ and finally, in Line 17 it updates $\hat{k}$ according to the new value of $\hat{m}$ in Line 15 which results in $\hat{k}=1$. Since $\hat{k}=1$, it means CPGR extracts only trivial bipartite cliques (shown in Figure 10(d)) which do not contribute to reduction in the number of edges in the restructured graph. Thus, it does not meet the condition in the while loop (Line 6) and CPGR terminates.
@@
CPGR then restructures the graph by adding two vertices, $z_{1}$, and $z_{2}$, corresponding to the two cliques $C_{1}$ and $C_{2}$, and adds the corresponding edges to form the tripartite graph. The edges in the given graph $G(U,W,E)$, that are not part of any $\delta$-clique are connected directly as shown in Figure 10(e).